TCO 8) Tomcat Company is planning to acquire a $250,000 machine to improve manufacturing efficiencies, thereby reducing annual cash operating costs (before taxes) by 80,000 for each of the next five years. The company estimated the weighted average cost of capital (WACC) is 8%. The machine will be depreciated using straight-line method over a five year life with no salvage value. Fritz is subjected to a combined 40% income tax rate. Required. A. What is the estimated net present value (NPV) of the proposed investment B. What is the payback period?
Answer #1

SOLUTION

Years 0 1 2 3 4 5
Machine Cost -250,000
Depreciation (250,000/5) 50,000 50,000 50,000 50,000 50,000
Tax saving in depreciation @ 40% (A) 20,000 20,000 20,000 20,000 20,000
Annual Benefit (B) 80,000 80,000 80,000 80,000 80,000
Net Cash Flow (A+B) -250,000 100,000 100,000 100,000 100,000 100,000
Payback Period -250,000 -150,000 -50,000

Payback Period = 2 + 50,000/100,000 = 2 + 0.5 = 2.50 years

Present value of cash flows = $100,000 * PVF (8%,5 years)

= $100,000 * 3.99271

= 399,271

Net present value = 399,271 - 250,000 =149,271